On April 1, Year 1, PPI purchased a forklift by paying $5,000 down and $5,000 at the beginning of each of the next 19 calendar quarters. What was the cost of the forklift for financial accounting purposes if the rate of interest agreed upon was 16% compounded quarterly?
Answer #1
The cost of the forklift will be the PV of the payments
discounted at the rate of 4% quarterly.
= 5000+5000*(1.04^19-1)/(0.04*1.04^19) = $          70,670
or
Alternatively, as an annuity due
= 5000*1.04*(1.04^20-1)/(0.04*1.04^20) = $          70,670